OpenStudy (ash2326):

@perurabo here you go The first jar contains 4 blue and 5 red marbles; the second basket contains 3 blue and 4 red marbles. One marble is randomly extracted from the first basket and put into the second. After that, a marble is extracted from the second basket. What is the probability that this marble is blue? A) 1/3 B) 15/36 C) 31/72 D) 4/9

OpenStudy (anonymous):

oh, ok, sorry

OpenStudy (ash2326):

No problem

OpenStudy (anonymous):

You know , this is kind of easy, but since people want things spelled out to them, I'll pass

OpenStudy (anonymous):

Among all possible scenarios there are two that suit us: I. A blue marble is put into the second basket and then a blue marble is extracted from the second basket. II. A red marble is put into the second basket and then a blue marble is extracted from the second basket. ======================================== A::The probability of the first scenario which is probability that a blue marble is taken from the first basket × probability that a blue marble is then extracted from the second basket = 4/9 × 1/2 = 4/18. B::The probability of the second scenario: probability that a red marble is taken from the first basket × probability that a blue marble is then extracted from the second basket = 5/9 × 3/8 = 15/72. =============================================== The probability of futher first or second scenario: 4/18 + 15/72 = ??? Add them and deduce the answer :)

OpenStudy (anonymous):

Either*

OpenStudy (anonymous):

so that gives 31/72 ?

OpenStudy (anonymous):

Exactly :3

OpenStudy (ash2326):

That's correct @Eyad Good :D

OpenStudy (anonymous):

oh, ok, thats great , ty

OpenStudy (anonymous):

Yw :)

OpenStudy (anonymous):

:)